OpenStudy (anonymous):

A machine fills jars with 9 ounces of jam. The machine is not exact and jars have a standard deviation of 0.05 ounce. In a lot of 1000 jars, how many jars would you expect to have between 8.9 and 9.1 ounces of jam? Give the exact answer. The probability a jar has between 8.9 and 9.1 ounces is at leas

OpenStudy (dumbcow):

assuming a normal distribution with mean 9 and std dev .05 9.1 - 9 = 0.1 8.9-9 =-0.1 0.1 is 2 deviations from mean -0.1 is -2 deviations from mean using z-table P(-2<x<2) = P(x<2) - P(x<-2) =0.9772 - 0.0228 = 0.9544 0.9544*1000 = 954.4 You would expect 954 jars to have between 8.9 and 9.1 oz of jam